determine slope of 3x + 5y =10
I got for y(0,2) and x(0, 10/3)
my answer was M=2/3.33
Is this correct, or where did I go wrong, please help. Thank you.

Answer by nerdybill(7384)

The two points you found:
y(0,2) and x(0, 10/3)
represents the y and x intercepts (where the line crosses the axis)
EXCEPT you wrote it wrong, it should be:
y-intercept=(0,2) and x-intercept=(10/3,0)
.
Then , the general formula to calculate the slope between two points is:
M = slope = rise/run = (Y2-Y1)/(X2-X1)
Plugging in our numbers:
M = (0 - 2)/(10/3-0)
M = (-2)/(10/3)
M = (-6)/10
M = -3/5
.